Two sides of the area measure 20 feet and two sides measure 10 feet. How many feet of fencing needed?
Answer by stanbon(75887)

Two sides of the area measure 20 feet and two sides measure 10 feet. How many feet of fencing needed?
----
Ans: 2[20+10] = 2*30 = 60 ft
-------------
Cheers,
Stan H.
------------------